OpenStudy (anonymous):

What about 7- |-23| + |-7|=n n=? :( Don't get it

OpenStudy (anonymous):

7 - 23 + 7 = -9

OpenStudy (anonymous):

Ok so 7-(23)+7 n=-9 the absolute value bars mean that the number inside has to be pos so, l-7l=7 and l7l=7

OpenStudy (anonymous):

right

OpenStudy (anonymous):

But if there is a negative outside the bars then you distribute it once you find the absolute value of the number.

OpenStudy (anonymous):

right - no confusion this time

OpenStudy (anonymous):

ok i get it now